OpenStudy (anonymous):

how do I factor a^2+ab+b^2?

OpenStudy (anonymous):

is it (a+b)(a+b)?

Directrix (directrix):

It will not factor over the set of real numbers. (a+b)(a+b) --> No.

OpenStudy (anonymous):

oh okay... thanks

Directrix (directrix):

(a+b)(a+b) = a^2 + 2ab + b^2

OpenStudy (anonymous):

how about a^3-b^3?

Directrix (directrix):

Yes. That is known as the difference of 2 cubes. It factors.

OpenStudy (anonymous):

what I mean is that how do i factor than one?

Directrix (directrix):

( a - b ) ( a² + ab + b² )

OpenStudy (anonymous):

don't you think that I should be (a-b)(a^2-b^2)?

Directrix (directrix):

Multiply it out and that will tell you that it is not. Now, I do agree with you that the expression looks as if it should factor as you indicated. It just doesn't.

Directrix (directrix):

Factoring Patterns: a^3 + b^3 = (a + b) (a^2 – ab + b^2) a^3 – b^3 = (a – b ) (a^2 + ab + b^2) -------- a^3 ± b^3 = (a [same sign] b)(a^2 + [opposite sign] ab + [always positive] b^2) I remember the formula in words as above. When you have time, read about factoring at the following link: http://www.purplemath.com/modules/specfact2.htm

OpenStudy (anonymous):

thanks..I understand it now.